### v1

# Upper Mid-Band Channel Measurements and Characterization at 6.75 GHz FR1(C) and 16.95 GHz FR3 in an Indoor Factory Scenario

Abstract
This paper presents detailed radio propagation measurements for an indoor factory (InF) environment at 6.75 GHz and 16.95 GHz using a 1 GHz bandwidth channel sounder. Conducted at the NYU MakerSpace in the NYU Tandon School of Engineering campus in Brooklyn, NY, USA, our measurement campaign characterizes a representative small factory with diverse machinery and open workspaces across 12 locations, comprising 5 line-of-sight (LOS) and 7 non-line-of-sight (NLOS) scenarios. Analysis using the close-in (CI) path loss model with a 1 m reference distance reveals path loss exponents (PLE) below 2 in LOS at 6.75 GHz and 16.95 GHz, while in NLOS PLE is similar to free-space propagation. The RMS delay spread (DS) decreases at higher frequencies with a clear frequency dependence. Similarly, RMS angular spread (AS) measurements show wider spreads in NLOS compared to LOS at both frequency bands, with a decreasing trend as frequency increases. These observations in a dense-scatterer environment demonstrate frequency-dependent behavior that deviate from existing industry-standard models. Our findings provide crucial insights into complex propagation mechanisms in factory environments, essential for designing robust industrial wireless networks at upper mid-band frequencies.
Index Terms:
FR3, FR1(C), 6G, InF, factory, upper mid-band, delay spread, angular spread, path loss, propagation, channel model
I Introduction
The increasing demand for higher data rates, lower latency, and enhanced reliability in wireless communication systems has driven the exploration of new frequency bands, such as FR1(C) (4-8 GHz) and FR3 (7-24 GHz). These upper mid-band frequencies offer a promising middle ground between coverage and capacity, providing wider bandwidth than the Sub-6 GHz spectrum while avoiding some of the propagation challenges associated with millimeter-wave (mmWave) frequencies. Regulatory agencies, such as ITU and FCC, have identified these bands as key candidates for future 5G and 6G deployments in urban and industrial environments [1, 2].
While significant research has focused on the higher mmWave and sub-terahertz (sub-THz) bands [3], there is still a limited understanding of the propagation characteristics in industrial environments, where wireless signals face distinct challenges. Unlike urban outdoor or indoor office spaces, factories often include diverse metallic surfaces, heavy machinery, and dynamic obstacles, which make propagation conditions more complex due to increased reflection, scattering, and diffraction effects. The propagation characteristics in these environments must be well understood to develop robust wireless systems for industrial applications [4, 5].
Previous studies have demonstrated the unique propagation behavior in industrial environments. For example, Rappaport et al. [6, 7] conducted pioneering measurements at 1.3 GHz in factory settings, reporting a path loss exponent (PLE) of 2.15, highlighting the need for dedicated channel models for factory environments. At higher frequencies, Schmieder et al. [8] found that at 28 GHz in industrial spaces, blockage significantly affects the LOS component, though it has a limited impact on overall channel gain.
Recent studies have explored various industrial environments across different frequency bands to better understand wireless channel characteristics. Vijayan et al. [9] conducted 5G channel measurements in a medical device manufacturing facility, focusing on the challenges posed by metallic structures. Using a simulated environment at 3.6 GHz, authors in [9] reported satisfactory channel gains despite the significant reflections caused by the metallic infrastructure, which increased multipath effects. In another study, Al-Samman et al. [10] performed wideband measurements at 107-109 GHz in manufacturing halls and test rooms. Their study revealed PLE ranging from 1.6 to 1.9 and a substantial increase in delay spread in NLoS conditions, emphasizing the need for optimized designs in such high-frequency scenarios.
Further research by Wang et al. [11] examined channel characteristics in micro-drilling and milling areas at 28 GHz. Their results indicated a PLE of 1.41 in LOS conditions and 2.81 in NLOS conditions. Additionally, they observed significant angular spread and delay spread, suggesting that propagation characteristics in industrial environments at mmWave frequencies are more complex due to dense multipath components (MPCs). Meanwhile, Lyczkowski et al. [12] studied the propagation behavior in automated guided vehicle lanes across frequencies from 2.1 GHz to 5.1 GHz. They reported notably small NLOS PLE values that increased with frequency, showing values ranging from 1.12 to 2.56. Their work highlights the variability of propagation characteristics in industrial environments as frequency increases, especially in the presence of obstacles like heavy machinery.
This work presents a detailed empirical study of channel measurements at 6.75 GHz and 16.95 GHz within the NYU MakerSpace factory environment, focusing on critical metrics such as path loss, delay spread, angular spread, and polarization. The remainder of this paper is organized as follows: Section II describes the channel sounder, the indoor factory environment, and the measurement procedures. Section III presents a comprehensive analysis of channel characteristics at 6.75 GHz and 16.95 GHz, including path loss modeling for co-polarized and cross-polarized antenna configurations, directional and omnidirectional path loss models, a comparison of delay and angular spreads across different environments and frequencies, and a comparison with the standardized values from 3GPP TR 38.901 [13]. Finally, Section IV summarizes the key findings of this paper.
II Channel Measurement in InF with a 1 GHz Bandwidth Channel Sounder
Figure 1: Floor plan of NYU MakerSpace factory environment showing three TX locations ($\bigstar$) and their associated RX positions (
$\bullet$
), with matching colors indicating TX-RX pairs.
II-A Channel Sounding System
A sliding correlation-based wideband channel sounder, detailed in [14, 15], was used for channel propagation measurements. The system generated a 500 Mcps PN sequence at baseband, upconverted to a 1 GHz wideband RF signal at 6.75 GHz for FR1(C) and further upconverted to 16.95 GHz for FR3 using a heterodyne architecture. The RX employed a slightly slower 499.9375 Mcps PN sequence, resulting in a time-dilated PDP due to the sliding correlation.
The unique feature of this sounder, as highlighted in [16], is its dual-band co-located design, which allows seamless switching between 6.75 GHz and 16.95 GHz by changing the IF signal cable to the appropriate RF module. Detailed specifications of the sounder system are provided in [17, 18, 16].
II-B Calibration of the Channel Sounder System
The calibration procedure ensures accurate power, delay, and direction capture of multipath components (MPC). The complete calibration comprises three phases:
II-B1 Linearity and Power Calibration
The linear range of operation and system gain is confirmed by measuring the transmit power using a Keysight N1913A power meter and the N8487A power sensor. The TX power is kept within the FCC’s 35 dBm licensed limit. Free-space path loss (FSPL) measurements at a distance beyond five times the far-field ($D_{f}$) of the horn antennas verify system gain and correct attenuator settings. A 4 m separation at 1.5 m TX and RX height ensures a single LOS path for accurate calibration, adhering to [19]. The linearity and power calibration are conducted twice daily at the start and end of the day to ensure accurate capture of MPCs throughout the day.
II-B2 Time Calibration
To ensure accurate multipath delay measurement, the two Rubidium (Rb) clocks at the TX and RX are first synchronized overnight using a physical cable. Each day, during system startup, Raspberry Pi computers connected to the Rb clocks initiate a Precision Time Protocol (PTP) synchronization to continually synchronize the two Rb clocks when the physical cable is disconnected for untethered measurements [20]. After linearity calibration, the LOS peak in the PDP is circular-shifted to match the free-space propagation delay for the 4 m calibration distance. After each location’s measurement, the TX and RX return to their calibration spots to remeasure the LOS path, to compensate for any minor drift in propagation delay accumulated during the long-duration measurement sessions.
Unlike previous campaigns, this setup continuously synchronizes the Rb clocks and PDPs using patent-pending methods [20], eliminating the need for post-processing or ray-tracing corrections to the time drift.
II-B3 Spatial Calibration
At each TX-RX location, antennas are mounted at 2.4 m (TX) and 1.5 m (RX) heights, using geographical North as the 0-degree reference for spatial coordination. Once calibration (system linearity, time, and space) is complete, the channel sounder is ready for propagation measurements [16].
II-C Indoor Factory Environment
[…]
During channel measurements, 3 TX locations and 12 RX locations were used, with the TX placed at a height of 2.4 m, and the RX at 1.5 m. Concrete pillars (represented as gray-filled circles in Fig. 1) were positioned throughout the space, impacting the MPCs propagation. Alongside the pillars, the various machines and material surfaces contributed to the complex scattering environment.
The TX-RX location pairs for the 6.75 GHz FR1(C) and 16.95 GHz FR3 measurements are listed in Table I, including 5 LOS and 7 NLOS TX-RX pairs, with separations ranging from 9 m to 38 m. At full power, our measurement system has a link margin of 155.6 dB at 6.75 GHz and 159.2 dB at 16.95 GHz[16, 21, 22]. However, to stay within FCC license requirements, a max EIRP of 31 dBm is used during field measurements to achieve a practical link margin of 142.7 dB at 6.75 GHz and 143.8 dB at 16.95 GHz [23], which results in no outages at any location.
TABLE I: TX-RX location pairs for factory InF measurements at 6.75 GHz in FR1(C) and 16.95 GHz in FR3 [16]
TX Loc Type RX T-R Separation (m)
TX1 LOS RX1 14.6
RX4 9
NLOS RX2 16.6
RX3 11.1
RX5 19
TX2 LOS RX1 33.3
RX3 24.9
NLOS RX2 38.2
RX4 24.6
TX3 LOS RX1 19.3
NLOS RX2 9.5
RX3 14
II-D Measurement Procedure
The measurements at the TX-RX locations follow a systematic approach as highlighted:
  • Rapid RX azimuth scans identify significant TX angles of departure (AODs) [16].
  • Detailed RX azimuthal sweeps in half-power beamwidth (HPBW) steps conducted for each AOD identified by the rapid scans following Table II.
  • Measurements in both co-polarized and cross-polarized antenna configurations.
  • Calibration before and after measurements at each location for temporal and spatial accuracy.
TABLE II: TX/RX elevation configurations for 360 azimuthal sweeps at fixed TX angles [16]
Sweep # TX elevation RX elevation
1 RX at boresight elevation, then swept 360 in azimuth plane in HPBW steps.
2 TX at boresight elevation. RX tilted down by one HPBW, then swept 360 in azimuth plane in HPBW steps.
3 RX tilted up by one HPBW, then swept 360 in azimuth plane in HPBW steps.
4 TX tilted down by one HPBW. RX at boresight elevation, then swept 360 in azimuth plane in HPBW steps.
5 TX tilted down by one HPBW. RX tilted down by one HPBW, then swept 360 in azimuth plane in HPBW steps.
HPBW: 30 @ 6.75 GHz; 15 @ 16.95 GHz.
III 6.75 and 16.95 GHz InF Radio Propagation Characteristics
III-A Path Loss
The measured path loss obtained from the directional and omnidirectional PDPs is fit to the CI model with a 1 m free space reference distance, $d_{0}$, to evaluate the path loss exponent (PLE or $n$).
[…]
$\text{FSPL}(f_{c},1m)$ $=32.4+20\log_{10}\left(\dfrac{f_{c}}{1\;\text{GHz}}\right),$
where FSPL$(f_{c},1\;\text{m})$ is obtained for carrier frequency $f_{c}$ GHz at 1 m, $n$ is the PLE, and $\chi_{\sigma}$ is large-scale shadow fading (zero-mean Gaussian r.v. with s.d. $\sigma^{CI}$ in dB) [24].
[…]
After synthesizing the omnidirectional PDPs using the method described in [25], we observe a lower PLE at 6.75 GHz, measuring 1.39 in LOS and 1.78 in NLOS, compared to 16.95 GHz, where the PLE values are 1.75 in LOS and 2.11 in NLOS. The low PLE at 6.75 GHz may be attributed to the structural characteristics of the MakerSpace, which includes plasterboard or wooden partitions with open sections at the top or bottom, facilitating easier penetration of lower-frequency waves. The relatively low PLE in NLOS at 6.75 GHz suggests a quasi-line-of-sight (QLOS) scenario, as defined in [26], where partial obstructions allow for some signal diffraction around obstacles, yielding a PLE less than 2, unlike fully obstructed NLOS conditions. Additionally, the absence of permanent walls contributes to the reception of scattered power from multiple directions at NLOS locations. The frequency comparison in Table III further highlights the trend of increasing PLE at higher frequencies.
3GPP does not include a CI model option for path loss in InF environments; instead, it uses the Floating Intercept (FI) model, as specified in Table 7.4.1-1: Path Loss Models in [13]. The FI model also varies the y-intercept as a fitting parameter and does not use a physical reference of the free-space propagation as in the CI model. Furthermore, 3GPP classifies the factory environment for TX heights above the surrounding clutter as sparse clutter-high height (SH) and dense clutter-high height (DH). Models for both classifications are shown in Table III.
TABLE III: Directional and Omnidirectional CI PL model (FI model for 3GPP) parameters for 6.75 GHz FR1(C) and 16.95 GHz FR3 InF campaigns, with comparison to 142 GHz campaign in [27]
Campaign Distance (m) Antenna HPBW (TX/RX) Directional path loss Omni path loss
LOS NLOS Best NLOS LOS NLOS
n $\sigma$ (dB) n $\sigma$ (dB) n $\sigma$ (dB) n $\sigma$ (dB) n $\sigma$ (dB)
6.75 GHz (This work) 9-38 (30/30) 1.64 1.44 2.22 3.88 2.92 7.69 1.39 1.86 1.78 2.46
6.75 GHz (3GPP) 10-600 $n$:2.15 $\alpha$:31.84 4.3 $n_{SH}$:2.3 $\alpha_{SH}$:32.4 $n_{DH}$:2.19 $\alpha_{DH}$:33.63 $SH$:5.9 $DH$:4
16.95 GHz (This work) 9-38 (15/15) 2.09 6.2 2.48 4.2 3.81 10.42 1.75 3.1 2.11 3.29
16.95 GHz (3GPP) 10-600 $n$:2.15 $\alpha$:31.84 4.3 $n_{SH}$:2.3 $\alpha_{SH}$:32.4 $n_{DH}$:2.19 $\alpha_{DH}$:33.63 $SH$:5.9 $DH$:4
142 GHz [27] 9-38 (8/8) 2.1 1.8 3.12 7.1 4.92 11.3 1.86 2.2 2.74 5.1
Notation: SH = Sparse clutter and High base station height (Tx or Rx elevated above the clutter), DH = Dense clutter and High base station height.
Figure 2: InF directional CI PL models and scatter plot for V-V polarization at: (a) 6.75 GHz FR1(C); (b) 16.95 GHz FR3. [T-R separation: 9–38 m]
Figure 3: InF omnidirectional CI PL models and scatter plot for V-V polarization at: (a) 6.75 GHz FR1(C); (b) 16.95 GHz FR3. [T-R separation: 9–38 m]
III-B RMS Delay Spread
The RMS DS provides a measure of the temporal spread of MPCs in the environment. A power threshold of 25 dB below the peak or 5 dB above the noise floor is used to evaluate the RMS DS. Table IV presents the RMS DS at the two frequencies, along with the 3GPP results from Table 7.5-6 Part-3: Channel Model Parameters for InF in [13]. The mean values are evaluated in log-scale and the expectations are calculated assuming underlying log-normal distribution for a direct comparison with 3GPP. NLOS RMS DS is clearly observed decreasing with frequency, LOS RMS DS is observed similar with slightly higher DS at 16.95 GHz, compared to 6.75 GHz. In LOS, the omnidirectional RMS DS is obtained as 14 ns at 6.75 GHz and 12.7 ns at 16.95 GHz, while in NLOS RMS DS is observed as 30.3 ns at 6.75 GHz and 29 ns at 16.95 GHz.
The RMS DS in the 3GPP model depends on the physical dimensions of the environment, particularly the volume (V) and enclosing surface area (S) considering the walls, ceiling and floor, rather than the frequency [13]. For the NYU MakerSpace floorplan shown in Fig. 1, the V is calculated as 3470 $\text{m}^{3}$ and the S as 1958 $\text{m}^{2}$, assuming a uniform height of 5 m.
TABLE IV: RMS Delay Spread Characteristics from 6.75 GHz to 142 GHz for LOS and NLOS Conditions.
Frequency (GHz) 6.75 6.75 (3GPP) 16.95 16.95 (3GPP) 142[27]
Dir RMS DS
LOS
$\mathbb{E}(\cdot)$ (ns)
11.2 13.7 1.8
NLOS
$\mathbb{E}(\cdot)$ (ns)
11.3 9.2 3.2
Omni RMS DS
LOS
$\mathbb{E}(\cdot)$ (ns)
14 26.84 12.7 26.84 5.5
NLOS
$\mathbb{E}(\cdot)$ (ns)
30.3 30.92 29.0 30.92 11
TABLE V: Angular Spread Characteristics measured by NYU WIRELESS at 6.75 GHz and 16.95 GHz in LOS and NLOS for the InF environment and Comparison to 3GPP Models [13]
Metric Condition 6.75 GHz 16.95 GHz
[…]
Omni RMS $\lg_{\text{ASA}}$ $=\log_{10}$($\text{ASA}/1^{\circ}$) $\mu_{\text{lgASA}}^{\text{LOS}}$ 1.25 1.62 23.84 46.56 1.05 1.55 14.13 40.86
$\sigma_{\text{lgASA}}^{\text{LOS}}$ 0.50 0.31 0.46 0.35
$\mu_{\text{lgASA}}^{\text{NLOS}}$ 1.80 1.72 66.91 58.21 1.61 1.72 50.42 58.21
$\sigma_{\text{lgASA}}^{\text{NLOS}}$ 0.21 0.3 0.43 0.3
Omni RMS $\lg_{\text{ASD}}$ =$\log_{10}$($\text{ASD}/1^{\circ}$) $\mu_{\text{lgASD}}^{\text{LOS}}$ 1.43 1.56 36.01 39.01 1.34 1.56 22.88 39.01
$\sigma_{\text{lgASD}}^{\text{LOS}}$ 0.50 0.25 0.20 0.25
$\mu_{\text{lgASD}}^{\text{NLOS}}$ 1.64 1.57 46.65 38.90 1.49 1.57 34.72 38.90
$\sigma_{\text{lgASD}}^{\text{NLOS}}$ 0.21 0.2 0.32 0.2
TABLE VI: Large scale spatio-temporal statistics [21]
Freq. TX RX Loc. TR Sep. Omni Abs. PL (V-V) Omni Abs. PL (V-H) Mean Dir. DS Omni DS Mean Lobe ASA Omni ASA Mean Lobe ASD Omni ASD Mean Lobe ZSA Omni ZSA Mean Lobe ZSD Omni ZSD
       [GHz] [m] [dB] [dB] [ns] [ns] [] [] [] [] [] [] [] []
       6.75 TX1 RX1 LOS 14.6 67.04 80.8 5.9 26.7 6.1 55.2 11.4 31.3 13.5 18.2 11.1 10.7
RX2 NLOS 16.6 69.81 87.44 8.8 32.7 7.4 50.4 11.0 26.5 7.6 14.8 4.5 11.3
RX3 NLOS 11.1 67.6 84.32 7.0 23.0 83.8 83.8 14.5 24.4 24.5 24.5 2.6 9.8
RX4 LOS 9.0 63.12 77.38 13.2 8.0 5.8 5.8 5.9 12.3 5.8 17.9 5.9 13.4
RX5 NLOS 19.0 71.38 83.26 5.6 27.3 124.0 124.0 18.3 28.2 1.1 7.6 2.9 9.2
TX2 RX1 LOS 33.3 72.02 89.56 3.0 14.6 7.9 62.1 13.7 41.6 9.3 20.4 1.8 8.9
RX2 NLOS 38.2 74.82 96.41 4.3 27.9 19.4 75.8 10.4 74.7 3.9 13.8 4.1 11.9
RX3 LOS 24.9 65.55 87.68 7.5 6.1 6.1 6.1 5.9 6.8 6.1 19.0 5.9 13.4
RX4 NLOS 24.6 79.06 96.65 13.5 40.5 47.7 47.7 48.2 48.2 12.6 12.6 1.7 5.7
TX3 RX1 LOS 19.3 65.52 84.35 10.2 18.7 15.2 15.2 15.9 133.3 2.4 13.4 6.1 11.3
RX2 NLOS 9.5 63.62 82.58 7.6 24.6 28.5 28.5 47.0 57.3 9.5 13.0 2.7 8.8
RX3 NLOS 14.0 69.85 87.58 4.8 38.6 26.6 78.4 72.0 75.9 13.1 13.3 2.3 7.0
       16.95 TX1 RX1 LOS 14.6 74.07 99.07 13.2 17.3 7.7 7.7 7.2 17.2 2.1 7.6 4.1 7.0
RX2 NLOS 16.6 78.09 104 11.3 56.1 7.5 30.7 9.5 18.6 5.4 9.6 3.9 6.4
RX3 NLOS 11.1 79.33 99.46 6.0 41.2 9.3 69.4 9.9 17.0 10.8 11.1 2.8 6.1
RX4 LOS 9.0 73.08 99.45 10.4 8.6 6.7 6.7 7.0 11.7 9.3 9.3 4.0 6.8
RX5 NLOS 19.0 88.54 116.5 3.3 2.8 5.8 5.8 8.8 10.5 8.9 11.5 2.8 5.9
TX2 RX1 LOS 33.3 82.63 105.7 8.5 22.0 6.8 71.6 9.6 33.6 7.1 11.8 2.0 5.5
RX2 NLOS 38.2 86.96 111.9 1.1 11.3 7.0 75.9 7.3 60.7 5.7 13.1 5.4 9.9
RX3 LOS 24.9 87.31 112.8 5.5 4.7 5.8 5.8 5.6 20.9 5.4 11.5 5.2 8.5
RX4 NLOS 24.6 90.65 100.7 9.0 43.8 19.5 64.7 14.5 65.6 2.8 6.0 3.6 5.3
TX3 RX1 LOS 19.3 77.97 98.77 10.9 13.9 8.0 8.0 6.9 35.3 2.2 7.8 1.6 5.1
RX2 NLOS 9.5 76.09 102.9 5.7 17.4 6.1 29.7 7.6 34.9 4.6 10.8 2.8 5.9
RX3 NLOS 14.0 82.15 105.6 3.9 50.5 14.1 106.0 15.3 58.5 5.9 7.7 3.8 6.9
III-C RMS Angular Spread
Figure 4: InF omnidirectional ASA: (a) 6.75 GHz; (b) 16.95 GHz. [T-R separation: 9–38 m]
The RMS AS shows the spatial dispersion of arrival or departure angles in azimuth and elevation planes. RMS AS is evaluated for a spatial lobe threshold of 10 dB below the strongest pointing direction, using procedures detailed in [22]. Table V summarizes the AS observed for both arrival angles at the RX (ASA) and departure angles at the TX (ASD). In LOS, the ASA is observed as 23.84 at 6.75 GHz and 14.13 at 16.95 GHz. Wide ASAs of 66.91 and 50.42 are observed in NLOS at 6.75 GHz and 16.95 GHz, respectively. 3GPP models a frequency-dependent LOS ASA with overestimated spread, along with a constant NLOS ASA, as outlined in Table 7.5-6 Part-3: Channel Model Parameters for InF and Table 7.5-11: ZSD and ZOD Offset Parameters for InF in [13]. ASAs are observed to decrease with frequency in both LOS and NLOS. Further, the ASDs are observed as 36.01 in LOS and 46.65 in NLOS at 6.75 GHz, while ASDs of 22.88 in LOS and 34.72 in NLOS are observed at 16.95 GHz. As shown in Fig. 4, the CDFs of empirical data align with these trends, while 3GPP predictions suggest a constant ASD of 39 in LOS and 38.9 in NLOS across frequencies.
Point data for large-scale spatio-temporal parameters corresponding to each of the TX-RX location pairs is summarized in Table VI. The format and procedure follows the detailed approach shown for the indoor hotspot environment in [21].
IV Conclusion
We presented an in-depth analysis of radio propagation behavior for InF at 6.75 GHz and 16.95 GHz, based on extensive measurements. 12 TX-RX locations were measured encompassing 5 LOS and 7 NLOS locations spanning 9–38 m. The path loss analysis revealed significant received power even in NLOS scenarios at 6.75 GHz. The PLEs were observed to be small with values below 2 in LOS and close to 2 in NLOS at both 6.75 and 16.95 GHz. The RMS DS was observed to decrease with increasing frequencies in both LOS and NLOS. The RMS AS also captured showed a decreasing trend with frequency and wide angular spreads were observed particularly in NLOS. Frequency-dependent trends for DS, and AS were observed in contrast to some of the constant modeling currently provided in 3GPP specifications.

### v2

@@
 […]
-Omni RMS $\lg_{\text{ASA}}$ $=\log_{10}$($\text{ASA}/1^{\circ}$) $\mu_{\text{lgASA}}^{\text{LOS}}$ 1.25 1.62 23.84 46.56 1.05 1.55 14.13 40.86
+Omni RMS $\lg_{\text{ASA}}$ $=\log_{10}$($\text{ASA}/1^{\circ}$) $\mu_{\text{lgASA}}^{\text{LOS}}$ 1.25 1.62 23.71 46.56 1.05 1.55 14.32 40.86
 $\sigma_{\text{lgASA}}^{\text{LOS}}$ 0.50 0.31 0.46 0.35
-$\mu_{\text{lgASA}}^{\text{NLOS}}$ 1.80 1.72 66.91 58.21 1.61 1.72 50.42 58.21
-$\sigma_{\text{lgASA}}^{\text{NLOS}}$ 0.21 0.3 0.43 0.3
-Omni RMS $\lg_{\text{ASD}}$ =$\log_{10}$($\text{ASD}/1^{\circ}$) $\mu_{\text{lgASD}}^{\text{LOS}}$ 1.43 1.56 36.01 39.01 1.34 1.56 22.88 39.01
+$\mu_{\text{lgASA}}^{\text{NLOS}}$ 1.80 1.72 66.38 58.21 1.61 1.72 50.40 58.21
+$\sigma_{\text{lgASA}}^{\text{NLOS}}$ 0.21 0.30 0.43 0.30
+Omni RMS $\lg_{\text{ASD}}$ =$\log_{10}$($\text{ASD}/1^{\circ}$) $\mu_{\text{lgASD}}^{\text{LOS}}$ 1.43 1.56 35.89 39.02 1.34 1.56 22.91 39.02
 $\sigma_{\text{lgASD}}^{\text{LOS}}$ 0.50 0.25 0.20 0.25
-$\mu_{\text{lgASD}}^{\text{NLOS}}$ 1.64 1.57 46.65 38.90 1.49 1.57 34.72 38.90
-$\sigma_{\text{lgASD}}^{\text{NLOS}}$ 0.21 0.2 0.32 0.2
+$\mu_{\text{lgASD}}^{\text{NLOS}}$ 1.64 1.57 45.93 38.90 1.49 1.57 34.77 38.90
+$\sigma_{\text{lgASD}}^{\text{NLOS}}$ 0.21 0.20 0.32 0.20
 TABLE VI: Large scale spatio-temporal statistics [21]
@@
 Table V summarizes the AS observed for both arrival angles at the RX (ASA) and departure angles at the TX (ASD).
-In LOS, the ASA is observed as 23.84 at 6.75 GHz and 14.13 at 16.95 GHz.
-Wide ASAs of 66.91 and 50.42 are observed in NLOS at 6.75 GHz and 16.95 GHz, respectively.
+In LOS, the ASA is observed as 23.71 at 6.75 GHz and 14.32 at 16.95 GHz.
+Wide ASAs of 66.38 and 50.40 are observed in NLOS at 6.75 GHz and 16.95 GHz, respectively.
 3GPP models a frequency-dependent LOS ASA with overestimated spread, along with a constant NLOS ASA, as outlined in Table 7.5-6 Part-3: Channel Model Parameters for InF and Table 7.5-11: ZSD and ZOD Offset Parameters for InF in [13].
 ASAs are observed to decrease with frequency in both LOS and NLOS.
-Further, the ASDs are observed as 36.01 in LOS and 46.65 in NLOS at 6.75 GHz, while ASDs of 22.88 in LOS and 34.72 in NLOS are observed at 16.95 GHz.
+Further, the ASDs are observed as 35.89 in LOS and 45.93 in NLOS at 6.75 GHz, while ASDs of 22.91 in LOS and 34.77 in NLOS are observed at 16.95 GHz.
 As shown in Fig.
@@
 Point data for large-scale spatio-temporal parameters corresponding to each of the TX-RX location pairs is summarized in Table VI.
-The format and procedure follows the detailed approach shown for the indoor hotspot environment in [21].
+The format and procedure follow the detailed approach shown for the indoor hotspot environment in [21].
 IV Conclusion